Jolene invests her savings in two bank accounts, one paying 5 percent and the other paying 8 percent simple interest per year. She puts twice as much in the lower-yielding account because it is less risky. Her annual interest is 6138 dollars. How much did she invest at each rate?
Amount invested at 5 percent interest is $
Amount invested at 8 percent interest is $

Answer by addingup(3677)

0.05(2x) + 0.08x = 6138
0.1x + 0.08x = 6138
0.18x = 6138
x = 34100 this is the amount invested at 8%
The amount invested at 5% is 2 times that amount = 34100*2 = 68200
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
At 5%: 68200 * 0.05 = 3410
+
At 8%: 34100 * 0.08 = 2728
-------------------------
Total . . . . . . . : 6138 Correct
.
Happy learning